Give an example of each of the following:


i. a function with a removable discontinuity


ii. a totally discontinuous function

1
Expert's answer
2022-03-02T17:07:10-0500

(i ) Example of a function with removable discontinuity

"f(x )= \begin{cases}
 \frac{x²-4}{x-2} &\text{if } x≠2\\
 1 &\text{if } x=0
\end{cases}"

Here f(x) is discontinuous at x=2. This discontinuity is removable discontinuity.

(ii) Example of a totally discontinuous function

"f(x) = \begin{cases}
 1 &\text{if } x &\ is &\ rational &\ number \\
 0 &\text{if } x&\ is &\ ir rational &\ number
\end{cases}" This function is totally discontinuous function.